first angle is twice the second angle, third angle is 20 degrees less than the second.
what is the measure of each angle

Answer by dfrazzetto(283)

Assuming a triangle:
a = 2b; thus b = a/2
c = b - 20
a + b + c = 180
put all in terms of one variable; we'll choose "b":
2b + b + b - 20 = 180
4b - 20 = 180
4b = 200
b = 50
a = 2b = 2(50) = 100
c = b - 20 = 50 - 20 = 30
a + b + c = 100 + 50 + 30 = 180